OpenStudy (anonymous):

Solve for m: m/5+3(m-1)/2=2(m-3)

OpenStudy (bahrom7893):

multiplying out by 10: 2m+15(m-1)=20(m-3)

OpenStudy (bahrom7893):

2m+15m-15=20m-60 -3m=-75 m=25 check the arithmetic